The mass of a high speed train is 4.5×105 kg, and it is traveling forward at a velocity of 8.3×101 m/s. Given that momentum equals mass times velocity, determine the values of m and n when the momentum of the train (in kg⋅m/s) is written in scientific notation.
1
Expert's answer
2020-10-30T04:03:36-0400

Mass "=4.5\times 10^8kg"

Velocity"=8.3\times 10^1m/s"

Momentum"=" Mass"\times" Velocity

In mathematical terms; "m\times v"

"\therefore" "P=[4.5\times 10^6Kg][8.3\times 10^1m/s]"

"=3.735\times 10^7kg/m/s"

In determination of the values of "m" and "n" from the momentum, the significant figure roles of multiplication apply, which are;

"3.7\times 10^7kg^{-1}m^{-1}s^{-1}"